Circular No. 7510

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2000 SV_{74} (LINEAR) An apparently asteroidal object reported by LINEAR on two nights

in September (first observation given below), and published on MPS 19881 under the designation 2000 SV_{74}, has been found to be cometary (diffuse with 16′′ coma and 20′′ tail at p.a. 150^{o}) by M. Tichý on CCD images taken on Oct. 19.8 UT with the 0.57-m f/5.2 reflector at Kleť.

2000 UT  α_{2000}  δ_{2000}  m_{2}
Sept.24.34197  1^{h}48^{m}07.^{s}62  +5^{o}13′14′′.9  17.8

Full astrometry (including observations by LINEAR on Sept. 5) and the following parabolic orbital elements are published on MPEC 2000-U07.

T = 2002 May 1.3060 TT  ω = 76.^{o}2417
Ω = 24.1918 }2000.0
q = 3.543939 AU  i = 75.1719
2000TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
Oct. 13 1^{h}34.^{m}75 + 5^{o}28.^{′}1 5.027 6.021 174.^{o}3 0.^{o}9 16.3
23 1 26.94 + 5 36.6 4.971 5.957 171.6 1.4 16.2
Nov. 2 1 19.16 + 5 46.9 4.949 5.894 160.4 3.2 16.2
12 1 11.76 + 6 00.2 4.961 5.831 148.9 5.0 16.1
22 1 05.08 + 6 17.5 5.002 5.768 137.4 6.7 16.1
Dec. 2 0 59.38 + 6 39.6 5.068 5.705 126.1 8.0 16.1
12 0 54.86 + 7 07.0 5.154 5.642 115.1 9.1 16.1
22 0 51.60 + 7 40.1 5.253 5.580 104.4 9.8 16.1

SUPERNOVA 2000dv IN UGC 4671 G. M. Hurst, Basingstoke, England, reports the independent discovery

of SN 2000dv (cf. IAUC 7509) by M. Armstrong, Rolvenden, on a CCD image taken on Oct. 17.136 UT (details as given for SN 2000dt on IAUC 7508). The magnitude for SN 2000dv was then 17.5, and the following position end figures were measured by Armstrong: 43^{s}.32, 13′′.3.

COMET C/1999 T1 (McNAUGHT-HARTLEY) Visual m_{1} estimates: Sept. 23.85 UT, 9.9 (A. Pearce, Nedlands, W.

Australia, 0.20-m reflector); Oct. 2.71, 9.6 (D. A. J. Seargent, The Entrance N.S.W., 25×100 binoculars); 18.83, 9.3 (Pearce).

2000 October 19  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green